# PRISM2: Unlocking Multi-Modal General Pathology AI with Clinical Dialogue

### v1

@@ 1 Introduction
The field of computational pathology has been transformed by the introduction of foundation models. Models such as Virchow2 [zimmermann2024virchow2], UNI2 [chen_general-purpose_2023] and H-optimus-1 [h-opt-1] are trained using millions of histopathology tiles spanning diverse tissue types, employing self-supervised objectives [oquab2023dinov2] to learn generalizable representations. The introduction of foundation models into the computational pathology modeling pipeline has improved performance, robustness, and data efficiency in a variety of tasks such as cancer detection, subtyping, and biomarker quantification [wang2022transformer, filiot2023scaling, chen_general-purpose_2023, vorontsov2024virchow].
Figure 1: PRISM2 uses language data derived from free-text clinical reports to align a visual representation combining multiple gigapixel WSIs with concepts relevant to pathology. For downstream task adaptation, whereas PRISM only produces an embedding vector of this aligned representation (here: ‘base’), PRISM2 uses a large language model to further refine it toward a diagnostic-tuned embedding (here: ‘diagnostic’). The language model thus serves as a diagnostic reasoning block that is also capable of generating diagnostic summary text and answering diagnostic questions. Questions with yes/no answers provide a simple and robust method of zero-shot prediction, though PRISM2 can also handle open-ended questions.
However, using these tile-level models at the level of WSI prediction still requires training an additional network to aggregate the tile embeddings across tens of thousands of tiles, which can overfit when training data is limited. A complete pathology foundation model must therefore operate at the level of WSIs. Early slide-level foundation models in pathology explored unsupervised training via masked autoencoding with dilated attention and weak supervision from clinical free text reports via contrastive alignment to the WSI representation, yielding improved slide-level representations [xu2024whole, ding2024multimodal, shaikovski2024prism]. The clinical report is a natural choice for learning a slide-level embedding as it contains a variety of clinically-relevant concepts which provide a diverse signal to prevent overfitting. While this direction is promising, these works remain constrained by limited dataset and model scale and are well positioned to benefit from recent progress in language modeling.
In this work, we present PRISM2 (Fig. 1), a slide-level foundation model that refines clinical-report-based supervision through dialogue, using a 4 billion parameter large language model (LLM). Beyond broadly aligning the WSI representation to a clinical report, we use question answering to provide fine-grained diagnostic supervision. PRISM2 is the largest multi-modal slide-level pathology foundation model to date, trained on the largest corpus of data containing 2.3 million WSIs from diverse institutions, corresponding to 685 thousand distinct clinical reports.
PRISM2 is trained using a two-stage approach. In the first stage, a perceiver-based [jaegle2021perceiver] image encoder coupled with a pretrained language encoder, BioGPT [luo2022biogpt], is trained with a contrastive objective. In the second stage, these embeddings are used as input to a Phi-3 Mini [abdin2024phi] language backbone to generate reports and answer questions, permitting flexible decoding into natural language. This design decomposes the learning problem, first capturing image representations and then fine-tuning the LLM to interpret them in a conversational framework.
PRISM2 is evaluated on a suite of diagnostic classification and biomarker prediction tasks. The choice of architecture and two-stage training approach yield two types of slide-level embeddings: the base embedding learned using the first stage of contrastive training and the diagnostic embedding based on the language model’s hidden state learned in the second stage. Zero-shot evaluation is performed using question answering, as well as contrastive similarity scoring. Overall, we show improved performance by using the language capabilities, while maintaining state-of-the-art performance in applications where language-based definitions are less well-defined, e.g. in the case of biomarkers. We also show the value of further finetuning the diagnostic embedding. Taken together, these results demonstrate the value of large-scale, multi-modal training.
The utility of PRISM2 is also investigated in two qualitative settings. First, a region of interest analysis is presented to investigate the accuracy and relevance of the highly attended regions of the WSI using expert clinical review. Second, we present examples of chat dialogue demonstrating the conversational capabilities of the model.
2 Related Work
2.1 Whole-slide multi-modal representation learning
Tile-level computational pathology models trained with paired text have previously been proposed [huang2023visual, lu2024visual], however, finding sufficient text to be paired with image crops is challenging. This is because clinical diagnostic workflows in practice are performed on one or more slides and lack fine-grained pairing to specific regions. Prov-Gigapath [xu2024whole] paired a whole-slide-level representation with clinical reports by first training a slide-level representation using a masked autoencoding objective on tile embeddings and then using a contrastive vision-language alignment loss with slide-level reports. Although this approach enabled the use of pathology reports, creating a sufficiently generalizable slide-level embedding in the first stage is challenging. PRISM [shaikovski2024prism] was the first work that addressed multi-modal pretraining on the specimen level, where each report was aligned with multiple slides to produce a single multi-modal embedding. The network was trained end-to-end using specimen-level reports and Virchow [vorontsov2024virchow] tile embeddings aggregated via a perceiver-based model [jaegle2021perceiver] in a single step training approach based on CoCa [yu2022coca]. Subsequently, TITAN [ding2024multimodal] has also introduced an approach using diagnostic reports. TITAN, building on CONCH embeddings [lu2024visual], uses three separate stages to train a WSI-level multi-modal model. The first stage is used to pretrain a region-level model using only embeddings and a self-supervised objective. In the second and third stages, the model is finetuned on regions with synthetic captions generated by PathChat [lu2024pathchat], followed by diagnostic reports, with each report paired to a single slide. Training on a single slide per report reduces computational complexity of the task but limits the applicability of the method to specimen-level tasks. PathAlign [ahmed2024pathalign] adopted the BLIP-2 [li2023blip] framework to pretrain a specimen-level aggregator by aligning its specimen-level representation to short strings capturing the most significant diagnostic findings.
Various works have also considered the use of non-text, paired modalities. TANGLE [jaume2024transcriptomics] aligns slide-level representations to corresponding genomic profile representations using contrastive alignment. THREADS [vaidya2025molecular] expands to three modalities by adding transcriptomic profiles. mSTAR [xu2024multimodal] uses both reports and genomic profiles for three-way contrastive alignment with a shared slide representation.
2.2 Whole-slide report generation
Approaches that aim to perform clinical report generation are sometimes, but not always, overlapping with approaches that aim to learn slide-level multi-modal representations. Notably, methods based on the CoCa [yu2022coca] framework are trained to produce a joint multi-modal representation and reconstruct the input caption at the same time, acquiring caption generation capabilities as a result. Both PRISM [shaikovski2024prism] and TITAN [ding2024multimodal], described above, use CoCa and so have report generation capabilities. While aligning a diagnostic report with WSIs produces a representation with diagnostically relevant features, exploring less complete captions like specific question-answer pairs may be trickier within the CoCa framework, as the reduced information content in these captions may also reduce the information content of the aligned image representations.
Some works use a separate training stage to enable a model to generate diagnostic findings given one or more slides. HistoGPT [tran2025histogpt] follows the Flamingo [alayrac2022flamingo] framework to align a Perceiver Resampler [jaegle2021perceiver] with a frozen BioGPT [luo2022biogpt] language model in its second training stage; the first stage is used to pretrain Perceiver Resampler on a classification task, likely to mitigate overfitting on the small training set used (15 thousand dermatology WSIs). PathAlign [ahmed2024pathalign], following the BLIP-2 [li2023blip] framework, finetunes the specimen-level aggregator with the frozen PaLM-2 S [anil2023palm2] language model to generate short strings with diagnostic findings. The model is trained on 350 thousand WSIs. PolyPath [ahmed2025polypath] finetunes a Gemini 1.5 Flash [geminiteam2024gemini15] language model on the same dataset in a single stage using a LLaVA-style [liu2023visual] framework. Both PathAlign and PolyPath use pre-computed tile embeddings extracted at 10$\times$ magnification. SlideChat [chen2025slidechat] trains on WSI and text data with CONCH [lu2024visual] as a tile encoder, LongNet [ding2023longnet] as a slide encoder, and Qwen2.5-7B Instruct [qwen2024qwen25] as an LLM in two stages, the first to do domain alignment, and the second to do instruction tuning. CPath-Omni [sun2024cpathomni] trains a tile encoder in a CLIP-style training stage using Virchow2 and CLIP models. They then use a multi-resolution tiling scheme along with a latent query attention pooler as a slide encoder and Qwen2.5-14B Instruct [qwen2024qwen25] as an LLM to train a unified mutli-modal patch and WSI-level model. ALPaCA [gao2025alpaca] also does a multi-resolution tile encoding scheme using CONCH, and uses a hybrid LongFormer [beltagy2020longformer] and a Gaussian Mixture Model adapter to connect the WSI modality to Llama 3.1 8B [meta2024llama3].
2.3 Whole-slide question answering
Several of these works explicitly focus on question answering capabilities as well. Beyond their captioning data, SlideChat trains on 176 thousand visual question answering pairs for 4,915 WSIs which is comprised of a mix of open-ended and multiple-choice question/answer (QA) pairs. CPath-Omni also trains on patch-level and WSI-level question answering data, for the latter they use 16 thousand QA pairs for 7,312 WSIs which are also a mix of open-ended and multiple-choice questions. Additionally, ALPaCA’s QA training data consists of 341 thousand QA pairs for 35,913 WSIs made up of true/false and multiple choice questions. A common theme in all these works is that each leveraged a general purpose LLM to generate QA pairs based on WSI reports/descriptions.
3 PRISM2
We present PRISM2, a multi-modal slide-level pathology foundation model designed to enable comprehensive analysis of WSIs through multiple interaction paradigms. Specifically, we sought to train a model with four key capabilities:

### v2

@@ 1 Introduction
Figure 1: Overview of PRISM2, a slide-level foundation model capable of producing generalizable embeddings as well as direct prediction with clinical dialogue. a The input to PRISM2 is one or more WSIs associated with a patient-case. These large tissue images are cropped into tiles which are subsequently processed with the Virchow2 foundation model. The sequence of tile embeddings is then aggregated via a slide encoder into a base image representation (“Base Embedding”). The base embedding is then used as input, along with text, to a large vision-language model which produces a diagnostic representation (“Diagnostic Embedding”) and text. The diagnostic embedding is tailored to diagnostic tasks, such as the detection and identification of cancers, precursors to cancers, and benign conditions. Base embeddings are more general and are well suited for transfer learning to tasks outside of the diagnosis-focused training distribution, such as biomarker and survival prediction. b Examples of the four types of direct dialogue-based prediction templates: report generation, open-ended QA, yes-no QA, and multiple-choice QA. Yes-no and multiple-choice QA in particular enable directly querying the model for quantifiable predictions with probabilities that can be calibrated.
The field of computational pathology has been transformed by the development of foundation models. Models such as Virchow2 [42], UNI2 [9] and H-optimus-1 [6] are trained using millions of histopathology tiles spanning diverse tissue types and employing self-supervised objectives [28] to learn generalizable representations. The introduction of foundation models into the computational pathology modeling pipeline has improved performance, robustness, and data efficiency in a variety of tasks such as cancer detection, subtyping, and biomarker quantification [35, 9, 33, 36, 7]. However, a representational gap still exists as most of the desired predictive tasks are at the level of the pathology case, which typically consists of one or more WSIs, and very few models provide generalizable slide-level embeddings. Instead, current workflows predominantly start from tile representations which are aggregated to create task-specific models via weakly supervised learning. The lack of a slide-level representation results in several shortcomings. For instance, without pretraining, aggregation models are more easily overfit and lack generalizability. They are also likely to require more training data and be less robust to nuisance variation.
The challenge of creating slide-level foundation models is in choosing a supervisory signal that is useful for compressing the long sequences of tile representations while still being sufficiently generalizable for the large number of possible downstream tasks. Work to date has considered image-only [39, 8], paired molecular [18, 17, 32], and clinical reports [2, 29, 13, 34] as sources for such signal. Given the consistent observation of the value of scale in foundation models for many different fields [19, 41], we prioritize signal that is abundantly available and therefore focus our slide-level model pretraining on clinical reports.
Past approaches leveraging clinical reports have suffered several shortcomings. First, many datasets are small, with insufficient diversity to learn generalizable representations. Second, unprocessed pathology report text itself is not enough to provide sufficient supervision. There is heterogeneity in pathology reporting protocols, and these are not always followed; information density can vary accordingly. Finally, many approaches have focused on contrastive objectives to align images and text which misses the opportunity to build on the capabilities of large language models.
Figure 2: Overview of PRISM2 training data. a The training data can be described in terms of patients, cases, specimens, blocks, and slides, as shown. Clinical reports are paired at the level of specimens. Consequently, PRISM2 is trained at the specimen level to predict diagnostic findings in the clinical reports. b The distribution of tissues represented in the training data. c The distribution of specimens and WSIs between MSK samples and those submitted for second opinion from diverse external sources around the world. d Distribution of high-level conditions in the training data. e Overview of how dialogue examples are generated from clinical reports to be used during PRISM2 training. Every processing step (arrow) uses GPT-4o. Clinical report rewrites are split into clinical history and pathology diagnosis. Diagnostic summaries are produced without demographic, molecular, and sectioning information. Diagnostic findings are first listed, and then converted into yes-no and open-ended question-answer pairs. As the yes-no question distribution is biased toward the findings mentioned in clinical reports, we mine additional complementary yes-no question-answer pairs by assuming that findings which are not mentioned are negative in the specimen. Multiple choice QA is derived directly from the reports, following the CAP reporting guidelines.
In this work, we present PRISM2 (Fig. 1, additional detail Extended Data Fig. 1), a slide-level foundation model capable of clinical-grade predictive performance. PRISM2 is the largest multimodal slide-level pathology foundation model to-date and is trained on the largest corpus of data, comprised of 685 thousand distinct clinical reports, corresponding to 2.3 million WSIs originating from diverse institutions and encompassing different tissue types (Fig. 2a-d and Methods 4.1). PRISM2 learns to aggregate the signal from tile embeddings and use it as input to a 4 billion parameter LLM. Several types of text data are created from the clinical report to provide sufficient and diverse training signal (see Fig. 2 and Methods 4.2 for details). This enables PRISM2 to be used for report generation and question-answering (see Extended Data Fig. 2 for examples). While PRISM2 is trained to replicate clinical dialogue, the goal of this supervision is not to deploy the model as a conversational agent for clinical practice but to use dialogue as a rich supervisory signal that encourages the learning of semantically grounded and transferable slide-level representations.
PRISM2 is trained using a two-stage approach (see Methods 4.3 and 4.4 for additional details). In the first stage, a perceiver-based [16] slide encoder, using Virchow2 [42] tile embeddings as input, is trained via a contrastive objective as well as an auto-regressive objective, similar to CoCa [40, 29]. Clinical text is used in both objectives: BioGPT [25] is used to create embeddings and Phi-3 Mini [1] is the language model for the contrastive and auto-regressive objectives, respectively. In the second stage, the image encoder is frozen and only the language model is fine-tuned using the auto-regressive objective. The choice of architecture and two-stage training approach yield two types of slide-level embeddings: the base embedding, defined as the output of the slide encoder, and the diagnostic embedding, defined as the language model’s hidden state. Empirically, we observe that the base embedding generalizes better to tasks not typically observed in the clinical report such as biomarker prediction and prognosis. Alternatively, the diagnostic prediction is tuned to tasks such as cancer detection and subtyping.
PRISM2 is evaluated on a suite of diagnostic classification, biomarker prediction, and survival tasks in several different manners aiming to capture the breadth of slide-level foundation model applications. Direct prediction, i.e. without further training of PRISM2, is performed using question answering. Linear-probe evaluation, i.e. with an additional linear predictive model, is used to measure how well the embeddings generalize to both in- and out-of-distribution tasks. In all cases, evaluation data is held out at the patient level from PRISM2 and Virchow2 training data. Overall, we show improved performance by using the language capabilities, while maintaining state-of-the-art performance in applications not typically represented in the clinical report, e.g. in the case of biomarkers. Notably, we show that, without further training, PRISM2 can achieve clinical-grade performance for cancer detection. We conclude with an assessment of forward looking goals, moving beyond detection and subtyping and demonstrate the first application of a slide-level foundation model for clinical report completion, following College of American Pathologists (CAP) reporting guidelines. Taken together, these results demonstrate the value of large-scale, multimodal training.
2 Results
2.1 Clinical dialogue training achieves diagnostic-grade performance without task-specific training
Figure 3: Direct prediction through dialogue with PRISM2. a Yes-no QA matches or exceeds the performance of existing clinical-grade products for the detection of invasive cancer as compared to Paige Prostate, Paige Breast, and Paige Breast Lymph Node (BLN), when evaluated on the corresponding product testing datasets. Yes-no QA on pan-cancer and rare variants exceeds the contrastive performance of competing approaches. After training a linear classifier using the diagnostic embedding as input on a large pan-cancer dataset, PRISM2 pan-cancer performance improves, and continues to outperform competing approaches with similar adaptation. b The size of the pan-cancer training set used for linear adaptation is similar to the training sets of the three clinical models; however, prostate (light blue), breast (dark blue), and BLN (yellow) tissues are a relatively small subset of the pan-cancer dataset. c Direct prediction performs well on TCGA cancer subtyping with multiple-choice QA. n indicates number of samples; ++ and - indicate number of positive and negative samples, respectively. Error bars show the 95% confidence interval computed with bootstrapping. * direct prediction result tied for first place (p < 0.05, permutation test).
Unlike tile-level computational pathology foundation models which are trained using a pretext task, PRISM2 is trained to generate clinically relevant text. As a result, for certain tasks, the model can have sufficiently high performance without further training, a setting we call direct prediction. PRISM2 is able to perform direct prediction via its question answering capabilities. To-date, direct prediction in the computational pathology literature has focused on the contrastive setting (e.g. [13, 29]), where candidate labels are compared with image embeddings and the nearest candidate label, measured by distance in embedding space, is selected. Using cancer detection and subtyping tasks for evaluation, we measure the direct prediction of PRISM2 using question-answering and PRISM and TITAN using contrastive comparison (see Sec. 4.5 for further methodological details on the evaluation protocols, Sec. 4.6 for further details on the comparison models and Sec. 4.7 for details on the evaluation datasets).
[…]
(continued) performance of the survival specialist model, followed by PRISM2 Base. c A large overall survival (OS) cohort was used to train the survival specialist model as well as the PRISM2 Survival model, by fine-tuning the slide encoder (base embedding). d PRISM2 base embeddings perform at least as well as those of any other model for biomarker detection, as evaluated via linear probing. e The number of times each model scored in the top 1, top 2, and top 3 across the 10 MSK and 7 TCGA biomarker tasks (measured with area under the curve (AUC) rounded to two decimal points). n indicates number of samples; n slide, n class, n event, and n task indicate number of WSIs, classes, recurrence or death events, and tasks, respectively. Error bars show the 95% confidence interval computed with bootstrapping. * linear probe result tied for first place (p < 0.05, permutation test).
First, we demonstrate direct prediction performance on tissue-specific cancer detection (Fig. 3a). We compare to specialist commercial models, specifically Paige Prostate, Paige Breast and Paige Breast Lymph Node (BLN). These models were developed with rigorous clinical validation and large, curated datasets, as summarized in Fig. 3b. Paige Prostate has FDA De Novo approval for clinical use in the US. Paige Breast Lymph Node has an FDA Breakthrough Device designation. All three products are CE-IVD and UKCA marked for clinical diagnostic use in Europe and the UK. Using yes-no question answering (Supplementary Table S1), PRISM2 matches the cancer detection performance of Paige Prostate and Paige Breast, and outperforms Paige BLN without further training (see Fig. 3a). For comparison, neither PRISM nor TITAN match the performance of the clinical-grade products with direct (contrastive) prediction and have a particularly notable performance gap on BLN. PRISM and TITAN are able to improve with linear adaptation (shown in the shaded region) but still trail PRISM2 performance.
Moving to pan-cancer prediction, again PRISM2 has the best performance overall (Fig 3a). While PRISM2 direct prediction is similar to linear pan-cancer prediction on prostate, breast, and BLN product testing datasets, linear adaptation shows some improvement on the pan-cancer test set which includes 9 common and 7 rare cancer types by tumor origin tissue. This suggests that for some cancer types, model adaptation by training with PRISM2 diagnostic embeddings can improve performance over direct prediction. Indeed for all three models, linear adaptation improves performance. Nevertheless, as with direct prediction, only PRISM2 matches or exceeds the performance of all three clinical-grade products with linear adaptation.
We also consider direct subtyping prediction using TCGA data [37, 22], specifically for breast, lung and kidney cancers (Fig 3c; see Supplementary Tab. S2 for multiple-choice question prompts). In these cases, the models largely have the same performance, except PRISM in the kidney task.
To calculate balanced accuracy (Fig 3a), both the direct and linear predictions use a probability threshold at 0.5, whereas, the clinical-grade products each have a carefully tuned prediction threshold. However, the threshold-free AUC performance (Extended Data Fig. 5a) yields the same conclusions: PRISM2 direct prediction can match or exceed the performance of clinical-grade products in cancer detection but there exists no probability threshold with which either PRISM or TITAN could do the same, even with linear probing trained on a large-scale pan-cancer dataset.
2.2 PRISM2 diagnostic embeddings are well suited to diagnostic pathology tasks
As noted above, as a result of model architecture and algorithmic training design choices, PRISM2 generates two types of embeddings: a base embedding and a diagnostic embedding. As is standard for foundation models, we expect these embeddings to generalize to many downstream tasks. To evaluate these representations, we gathered several datasets to assess different models’ ability to identify cancers, precursors, and benign conditions. Specifically, we gather pan-cancer, pan-cancer (rare), breast, and GI detection tasks from MSK, and BRCA, NSCLC, and RCC subtyping from TCGA [37, 22]. The large MSK datasets present challenging real-world pathology tasks. The pan-cancer task involves the detection of invasive cancer across 16 cancer tissues of origin (7 of which comprise the “rare” cancers subset). The breast and GI tasks involve the detection of cancers, precursors to cancer, and benign conditions (see Extended Data Fig. PRISM2: Unlocking Multi-Modal General Pathology AI with Clinical Dialogue for detailed per-condition results). The TCGA tasks involve cancer subtyping in breast, lung, and kidney. Additional dataset details can be found in Sec. 4.7.
As comparison slide-level embeddings, we evaluate PRISM [29], TITAN [13], COBRA [20], Prov-GigaPath [38], and a tile-level baseline using only the mean of the Virchow2 [42] embeddings. The mean tile embedding is an important baseline as it represents a naive aggregation of the tile embeddings. In all cases, the slide-embedding is used as input to a linear model. For further details regarding the comparison methodology and models, refer to Sec. 4.5 and 4.6, respectively.
Fig. 4a presents performance on standard clinical tasks. In all settings, both PRISM2 base and diagnostic embeddings perform as well as or better than other embeddings. For pan-cancer detection in particular, PRISM2 diagnostic embeddings (0.976 AUC on all cancers) substantially outperform both PRISM2 Base embeddings (0.956 AUC) and all other embeddings (0.947 AUC with PRISM, 0.931 AUC with TITAN). Furthermore, the performance drop on the rare cancers subset is modest (0.967 AUC to 0.957 AUC). In breast, PRISM2 and PRISM embeddings outperform others. In GI, we observe a marked performance drop for TITAN on precursors to cancer and benign conditions (Extended Data Fig. PRISM2: Unlocking Multi-Modal General Pathology AI with Clinical Dialogue), whereas PRISM2 performance remains high across all conditions. Interestingly, although the COBRA slide-level model uses Virchow2 tile embeddings, its embeddings often under-perform the baseline method of taking a mean of the Virchow2 tile embeddings, suggesting the importance of the slide-level training stage.
2.3 PRISM2 survival embeddings outperform a survival specialist model
Accurate prognostic prediction is of major interest for the computational pathology field but generally understood to be very challenging both because of lack of data as well as uncertain connections between morphology and patient outcome [31, 30]. We assess the benefit of pretraining with report data to survival prediction tasks as well as the value of additional finetuning using a large survival dataset. To do so, we gather a dataset of over 225 thousand cases with overall survival (OS) information, across nearly 100 thousand patients (Fig. 4c). Using this data, we fine-tune the PRISM2 slide encoder as well as train a survival specialist model from scratch (see Sec. 4.4 for additional detail). We refer to the embeddings of the finetuned model as survival embeddings. We select colorectal cancer (CRC) recurrence-free survival RFS prediction with MSK data and 13 disease-specific survival DSS prediction tasks with TCGA data for evaluation (Sec. 4.7 for dataset details).
The results of the survival analysis are presented in Fig. 4b. PRISM2 survival embeddings exceed the performance of the survival specialist model, especially for MSK CRC RFS (0.809 vs 0.773 C-Index). Without survival fine-tuning, PRISM2 base embeddings outperform all other model embeddings for MSK CRC RFS and match the Virchow2 Mean baseline (mean of Virchow2 tile embeddings) on TCGA DSS. Unsurprisingly, the base embeddings slightly outperform the diagnostic embeddings, as survival information was not present in the dialogue training data. See Extended Data Fig. 8 for detailed per-task results.
2.4 PRISM2 base embeddings generalize to biomarker prediction
Biomarker prediction using hematoxylin and eosin (H&E) WSIs is another task that, like survival prediction, could extend the applications of computational pathology. Early results have shown great promise in this area [36, 7]. We select two sets of biomarkers, from MSK and TCGA to evaluate prediction performance using embeddings (Sec. 4.7 for dataset details). Fig. 4d,e (and Extended Data Fig. 9) presents the results. Similarly to survival prediction, PRISM2 base embeddings tend to outperform diagnostic embeddings. While the average AUCs on MSK TCGA tasks are the highest for PRISM2 base embeddings, as compared to diagnostic embeddings, PRISM, TITAN, Virchow2 Mean (mean of the tile embeddings), COBRA, and GigaPath, the next-best performing models are close: 0.846 (vs 0.854) AUC with COBRA on MSK and 0.781 (vs 0.784) AUC on TCGA with TITAN. Similarly, looking at Fig. 4e, PRISM2 embeddings (either base or diagnostic) score the most often among the top 1, top 2, and top 3 best-performing models per biomarker prediction task but only by a small margin (Fig. 4e). Overall, we can conclude that PRISM2 embeddings are, on average, at least as good as those of other models for biomarker prediction.
2.5 Dialogue unlocks pathology report completion
[…]
As a final demonstration of the value of direct prediction with dialogue, we aim to move beyond benchmark tasks, and evaluate PRISM2’s ability to complete routine pathology reports, following CAP reporting guidelines. We are the first to explore detailed pathology report completion directly with a pathology foundation model, without further training, and aim to provide a perspective on how a dialogue-based system could be used in real clinical practice.
We demonstrate this capability with the CAP report template for biopsies with invasive carcinoma of the breast [4] (see Sec. 4.5 for details on the methodology and evaluation and Sec. 4.7 for dataset details). The questions associated with report fields and detailed performance metrics are shown in Fig. 5. The fully shaded bars show the adjusted mean recall (AMR) achieved directly by question-answering (“observed AMR”), whereas the lightly shaded extensions to these bars show the maximum AMR achievable by calibrating the probabilities of the responses (“max AMR”). We can see that histologic type, ductal carcinoma in situ (DCIS) presence, DCIS architectural pattern, and lymphovascular invasion fields benefit substantially from calibration. On the other hand, tubule formation score, DCIS nuclear grade, necrosis, and microcalcifications benefit less. Overall, the results demonstrate that PRISM2 has modest to good understanding of the diagnostic concepts discussed but in some cases, the model’s responses cannot be used without calibration.
Selecting the correct histologic type among 20 options is a challenging task. Only fourteen options were represented in the evaluation data, with eight options (B, C, D, E, I, J, L, P) having at least ten examples (see ‘n’ and ‘+’ in Fig 5). PRISM2 achieved an AMR of 0.519 (0.731 with calibration) across these options. Mucinous carcinoma (I) and adenoid cystic carcinoma (P) are the only options where the observed AMR matches the max AMR (0.994 and 0.986 observed AMR, respectively); however, the observed AMR for IDC of no special type (B) and ILC (D) is low (0.557 and 0.276, respectively). From the TCGA BRCA results in Fig. 3c and the breast IDC and ILC results in Extended Data Fig. PRISM2: Unlocking Multi-Modal General Pathology AI with Clinical Dialoguef, we know that these indications can be well identified by the model. It appears that correctly identifying these options becomes difficult when including all six ductal carcinoma variant options (B, C, I, J, L, P), as well as the ductal and lobular mixed type (E). Indeed, a stratified analysis of histologic type classification reveals that IDC vs. ILC performance is good (0.933 AMR) and does not benefit from calibration. On the other hand, when pooling rare histologic subtypes as prognostic of high risk or low risk, calibration is required. This serves as an example of a downstream task that could be developed from the concepts presented in this clinical report if validation (or training and validation) data are available.
2.6 The diagnostic summary highlights key image regions
Finally, connecting predictions to morphologic features can be an important tool for verifying model predictions. To this end, we selected two samples for detailed pathologic review. We qualitatively evaluate the generated text and compare with attention maps, the sum of all the scores per tile across attention heads and queries multiplied by the $\ell_{2}$-norms of the corresponding values. As shown in Extended Data Fig. 3 and 4, PRISM2 produces verbose descriptions of specimens. These descriptions remain consistent when looking only at the cropped regions that the model attends to the most. The model’s attention maps, taken from the slide encoder’s cross-attention layer, are shown for a single core of a WSI and each of its top regions of interest.
3 Discussion
A complete pathology foundation model encodes the full range of relevant pathologic semantics not only within tiles, but also across WSIs. This minimizes the training and data required to adapt it to downstream tasks, while enabling generalization to diverse tasks like rare cancer detection and prognostic prediction. In this work, we achieve these aims. PRISM2 base embeddings outperform those of other slide-level models on diagnostic, prognostic, and biomarker tasks. PRISM2 diagnostic embeddings, which benefit from a language model trained on diagnostic dialogue, further improve performance on diagnostic tasks. Indeed, PRISM2 is the only model to match or exceed the performance of all tested clinical-grade products in cancer detection and does so with direct question answering without further training or adaptation. Finally, PRISM2 survival embeddings, trained on an overall survival cohort of unprecedented scale, containing nearly 100 thousand patients, further outperform the base embeddings on prognostic tasks including RFS and DSS.
Building on a language model was critical for unlocking PRISM2 direct prediction performance. Natural language provides a straightforward manner in which to frame clinical tasks, even when the positive or negative cases are hard to define. For example, contrastive direct prediction for pan-cancer detection requires all possible cancer-positive phrases to be contrasted with all possible non-cancer phrases, including benign conditions across all tissues (see Supplementary Tab. S3). Conversely, the same yes-no question (e.g. “Is malignant tumor present in this specimen?”) can be used across all pan-cancer data. This consideration may be less important when the number of candidate labels is small, as is often the case for subtyping. However, we note that we do not observe contrastive approaches outperforming PRISM2 direct prediction in these cases, thereby indicating that PRISM2 allows for general QA without sacrificing performance on more narrowly defined tasks. Moving to linear classification removes this consideration, but PRISM2 still outperforms competing models.
Optimization details were also important for achieving generalizability. We freeze the slide encoder early in training before reaching the lowest contrastive validation loss. We find this achieves a better balance between performance on diagnostic tasks and transferability to biomarker and prognostic tasks, in line with other work that also has observed that earlier checkpoints may be more generalizable [14]. We then rely on the vision-language model to refine the representation into one specialized for diagnosis. We have observed freezing the slide encoder prevents the vision-language model from overfitting and ultimately improves performance.
Although PRISM2 is capable of dialogue, accurate question answering is prioritized over robust chat capabilities. While abilities beyond our training set, like answering compound questions, are inherited from the Phi-3-Mini initialization, they are not used for downstream tasks. Instead, training the model to answer yes-no and multiple-choice questions, alongside open-ended questions and more general report generation provides us with three ways in which the vision-language model of PRISM2 can be adapted to downstream tasks: (1) by uncalibrated direct prediction, (2) by direct prediction of probabilities calibrated on a validation set, and (3) by further training using base or diagnostic embeddings. The first setting allows direct prediction of conditions without further training. Additionally, instead of directly using the PRISM2 responses, we can quantify the probability of response to a specific yes-no and multiple-choice question. Finally, the diagnostic embeddings are tuned for diagnostic pathology tasks such as the detection and identification of cancers, precursors to cancer, and benign conditions and can be adapated when additional training and validation data are available.
There are still open questions regarding directions of possible improvement for slide-level foundation models. Because the perceiver-based slide encoder relies on cross-attention rather than self-attention, we did not include position encoding. Also, although Virchow2 can encode tiles at many magnifications, PRISM2 only used embeddings at 0.5 microns-per-pixel (mpp). Moving to a different aggregator architecture with position encoding and mixed magnification input may improve performance on some tasks that benefit from greater spatial context, like cancer subtyping, and may enable measurement tasks, like mitotic event counting. As evidenced by the modest improvement in biomarker performance compared to baseline models, as well as the large improvement in survival prediction performance gained by fine-tuning on OS data, we posit that further improvement in prognostic and biomarker tasks will require including them in the model pre-training phase. Finally, while it appears that full diagnostic report completion is possible, it remains challenging. A complete study of report completion would evaluate inter-pathologist variability. Grading and subtyping are particularly subjective and may be subject to biases.
Work in foundation models in computational pathology has moved from the tile level to the slide level [29, 13, 34, 39, 18, 17, 32, 2]. PRISM2 extends the scale of this effort and is the first to achieve clinical-grade cancer detection performance. It is the largest model (4.6 billion parameters) trained on the most comprehensive dataset to date with nearly 700 thousand specimens and clinical reports, comprising over 2.3 million slides and 14M question-answer pairs. Besides the scale, we demonstrate through the inclusion of a large vision-language model that the image-report alignment methods to date can be viewed as a base for further foundation model development, and model performance can be further improved by borrowing from the large language model (LLM) domain. Finally, we find that diagnostic performance in pathology is beginning to saturate and that further improvements in biomarker and prognostic prediction, driven by the inclusion of this data in foundation model pre-training, may be the new frontier for foundation models in pathology.
4 Methods
4.1 Training Data
Institutional review board review was not applicable for the research described in this study. This research study was conducted retrospectively from deidentified data licensed to Paige.AI, Inc. from MSK. The data used in this study were all collected originally for clinical use by MSK in the practice setting and are therefore considered secondary data. Only data previously deidentified by MSK were utilized in the analysis, and unique patient identifiers were completely removed from the analytical dataset. To the best of our knowledge, MSK has not transferred any data for which the applicable patient has not consented to or otherwise agreed to MSK’s Notice of Privacy Practices or a substantially similar notice, waiver or consent.
[…]
WSIs are processed using a fully-convolutional network to perform tissue segmentation and tiled at $20\times$ magnification with a size of $224\times 224$ pixels, discarding any tiles that are less than 65% tissue. The resulting tiles are then processed by Virchow2 [42] to produce tile embeddings, specifically the 1280-dimensional class token output of the model. During training, up to 100K tiles are loaded per specimen. In cases where specimens have more than 100K tiles, WSIs in the specimen are dropped until the criterion is met. This restriction does not apply during inference.
Raw report data is delivered at the case-level, but reporting is often done at the specimen-level which requires extracting and matching each report to the proper set of WSIs for that specimen. The reports are a mix of synoptic worksheets and free-text and are stripped of boilerplate content like headers/footers, disclaimers, and repetitious formatting characters.
4.2 Creating clinical dialogue for training
To create high-quality text data for training, we apply GPT-4o [27] with various prompting strategies to the report data (see Fig. 2 for an example workflow). We create four classes of processed text tasks to be used in training: